# Spin Theory Based on the Extended Least Action Principle and Information Metrics: Quantization, Entanglement, and Bell Test With Time Delay

### v1

@@ Abstract
A theory of electron spin is developed here based on the extended least action principle and assumptions of intrinsic angular momentum of an electron with random orientations. By incorporating appropriate relative entropy for the random orientations of intrinsic angular momentum in the extended least action principle, the theory recovers the quantum formulation of electron spin. The two-level quantization of spin measurement is a natural mathematical consequence instead of a postulate. The formulation of measurement probability when a second Stern-Gerlach apparatus is rotated relative to the first Stern-Gerlach apparatus, and the Schrödinger-Pauli equation, are also derived successfully. Furthermore, we provide an intuitive physical model and formulation to explain the entanglement phenomenon between two electron spins. In this model, spin entanglement is the consequence of correlation between the random orientations of the intrinsic angular momenta of the two electrons. Since the orientation is an intrinsic local property of electron, the correlation of orientations can be preserved even when the two electrons are remotely separated. Such a correlation can be manifested without causal effect. Owing to this orientation correlation, the Bell-CHSH inequality is shown to be violated in a Bell test. The standard quantum theory of electron spin can be considered as an ideal approximation of the present theory when certain conditions are taken to the limits. To test the difference between the present theory and the standard quantum theory, we propose two potential experiments. First, the observation of spin quantization depends on the its interactions with the measuring magnetic field; Second, in a typical Bell test that confirms the violation of Bell-CHSH inequality, the theory suggests that by adding a sufficiently large time delay before Bob’s measurement, the Bell-CHSH inequality can become non-violated.
@@ I Introduction
With respect to spin entanglement between two electron spins, we will show that the root cause of entanglement is the correlation between the random variables, the orientations of the intrinsic angular momenta, of the two electrons. The correlation can be established during preparation. Crucially, since the orientation of the angular momentum is an intrinsic local property of the electron, the correlation can be maintained even though the two electrons are space-like separated and without interaction. There is no “spooky action in the remote” in the EPR experiment. Mathematically, we give an equivalent formulation of the four Bell states using probability density function instead of wave function. Based on the probability density function, Bell-CHSH inequality is proved to be violated in our theory due to the same root cause of entanglement. Since the correlation of the orientations of the angular momenta is preserved even when the two electrons are separated, the joint probability for the measurement outcome cannot be factorized into a product of two individual terms - a requirement for the Bell inequality to hold.
@@
The standard quantum theory of electron spin can be considered as an idealized approximation of the present theory when certain conditions are taken to limits. When these conditions are not taken to the limits, the theory predicts results that are different from standard quantum mechanics. Two potential experiments are proposed to test the difference. In the first experiment, the Stern-Gerlach apparatus is configured with a sufficiently weak gradient of the inhomogeneous magnetic field along the $z$-axis; We expect the electron detector screen will exhibit continuous distribution of displacements along the $z$-axis around two areas, instead of only two discrete lines. The second experiment is to modify a typical Bell test experiment that has already confirmed the violation of Bell-CSHS inequality. Here, we purposely let Alice and Bob perform their measurements at different times on each entangled pair of electrons. While Alice performs her measurement at time $t_{a}$, Bob purposely delays his measurement at time $t_{b}=t_{a}+\Delta t$. The present theory predicts that when $\Delta t$ is chosen properly, the Bell-CSHS inequality will not be violated.
@@
The rest of the paper is organized as follows. In Section II, we review the extended least action principle and its underlying assumptions. Detailed calculation showing how the basic quantum theory is derived from the principle is provided in Appendix A. Section III is devoted to developing the spin model, where the extended least action principle is incorporated with Tsallis relative entropy, resulting in a probability density function for the random orientation. The probability density function becomes two Dirac delta functions when the order of relative entropy approaches a limit, corresponding to the quantization of measurement outcomes. The theory for rotation of the spin is developed in this section as well. In Section IV, the Schrödinger-Pauli equation is derived by combining the effects of random translational fluctuations and random rotational fluctuations when applying the extended least action principle. Section V develops the theory for spin entanglement based on the correlations between the random orientations of intrinsic angular momenta, which allows us to prove that Bell inequality is violated with this spin theory. To test the difference between our theory and the standard quantum mechanics, we propose two potential experiments in Section VI. Lastly, the conceptual implications and the limitations of the present theory are discussed in Section VII.
@@ III.2 Discreteness of Spin Measurement
When an inhomogeneous field $B_{z}$ is applied, $m$ must be a number larger than zero. This means $m$ increases when the spatial gradient of $B_{z}$ is non-zero. We can assume further that $m$ monotonically increases as the electron travels along the inhomogeneous field $B_{z}$. This is intuitively explained as follows. The key is not to consider the electron as an idealized point particle. Instead it has non-zero size and is distributed with a spatial volume. We do not assume it is a rigid rotating body. Instead, it can be divided into many small mass elements $\delta m$. Each mass element is also a charge element $\delta e$, and experiences Larmor precession. Note that in Larmor precession, the angular frequency is $\omega=\delta eB_{z}/\delta m$. In a homogeneous magnetic field, $\omega$ is a constant for each charge element and independent of the angle $\theta$. In this case, the probability density function $p_{m}(\theta)$ should remain unchanged for the whole electron. But in an inhomogeneous field $B_{z}$, because the electron has spatial size, each charge element will experience a different magnitude of magnetic field. Consequently, different parts of the electron will have different angular frequencies. The upper parts of the electron along the direction of $B_{z}$ tend to precess faster than the lower parts of the electron. Effectively, the orientation of overall intrinsic angular momentum is “pulled” closer to align with the $z$-axis. This corresponds to choosing a larger value of $m$.
Although the above physical interpretation is intuitive, it is nevertheless an assumption and need to be called out explicitly as
@@ V.2 Bell States
In the case that the initial orientations of angular momentum of electron $A$ are uniformly distributed for all $\theta\in[0,\pi]$, one has $\varrho(\Theta^{+}_{A})=\varrho(\Theta^{-}_{A})=1/2$, (60) and (62) become
@@
On the other hand, if he correlations between the electron pair are such that $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ along the $z$-axis, and $\theta_{A}^{y}=\theta^{y}_{B}$ along the $y$-axis, we must also have
@@
Similarly, the combination of (63) and (65) is equivalent to the Bell state
@@ VI Possible Experiments to Confirm the Spin Theory
The spin theory presented here so far recovers many of the results from standard quantum mechanics. In fact, the standard quantum theory of electron spin can be considered as an ideal approximation of the present theory when certain conditions are taken to the limits. When these conditions are not taken to the limits, the present theory will predict some results different from standard quantum mechanics. In this section, we will give two possible experiments to confirm the difference.
@@ VII.1 Conceptual Implications
We see the spin model and theory presented here gives an intuitive physical picture that may help to consolidate several quantum physics interpretations. In addition, our theory supports the idea that measurement outcome is context dependent. Measurement outcome for the spin of an electron depends on the measurement context, that is, the direction of magnetic field used in the Stern-Gerlach apparatus. The two level discreteness of measurement outcome is the result of a sufficiently strong magnetic field. For an electron with initially completely random orientation of intrinsic angular momentum, we can measure the spin along any direction but always get spin-up and spin-down with equal probability. There is no predefined spin property without specifying an experimental context.
@@ VII.2 Limitations
Secondly, assumption 4 suggests that when the interaction between the electron and an inhomogeneous magnetic field in a Stern-Gerlach apparatus takes place, the order of relative entropy monotonically increases, so that the direction of intrinsic angular momentum tends to align with the direction of the magnetic field. We give an intuitive explanation by taking into consideration that the electron is not an idealized point particle. A more rigorous theory to confirm this assumption is needed. The spin theory based on stochastic mechanics Beyer gives similar results. However, stochastic mechanics relies on velocity fields which are non-local in nature, which makes stochastic mechanics a less favorable theory.
@@ VII.3 Conclusions
From the extended least action principle, and with additional assumptions on the intrinsic angular momentum of an electron, particularly its random orientations, we are able to develop a theory that recovers the quantum formulation for electron spin. The key component of the theoretical framework here is the introduction of relative entropy for both translational and rotational random fluctuations. The two-level quantization of spin measurement outcomes is no longer a postulate. Instead, it is a mathematical consequence based on Assumption 4. We also obtain the same formulation as standard quantum mechanics when the direction of magnetic field of a second Stern-Gerlach apparatus is rotated with an angle from the direction of magnetic field of the first Stern-Gerlach apparatus. Recursively applying the extended least action principle, we have derived the Schrödinger-Pauli equation.
@@
The standard quantum theory of electron spin can be considered as an ideal approximation of the present theory when certain conditions are taken to the limits. When these conditions are not taken to the limits, the present theory predicts results that are different from standard quantum mechanics. For instance, if the Stern-Gerlach apparatus is set up with a sufficiently weak gradient of the magnetic field along the $z$-axis, we expect the electron detector screen will exhibit continuous distribution of displacements along the $z$-axis, instead of only two discrete lines. Another more interesting experiment proposed is to add a delay before Bob’s measurement in the typical Bell test experiment, which could result in the non-violation of the Bell-CHSH inequality. The second experiment shows the effect of time dependency in measurements involving spin entanglement. If confirmed by experiment, such an effect can impose a limitation of the application of spin entanglement in quantum computing and quantum information.
The interplay of the extended least action principle and the information metrics for vacuum fluctuations prove to be very fruitful. We have shown the theoretical framework based on the extended least action principle has successfully recovered the non-relativistic quantum mechanics Yang2023 , the relativistic quantum scalar field theory Yang2024 , and in the paper, the quantum theory for electron spin. We believe the results in this paper are more interesting and conceptually more important because we can explain the entanglement phenomenon with an intuitive physical picture, and predict new results different from standard quantum mechanics that can be potentially tested in the proposed experiments.

### v2

@@ Abstract
A theory of electron spin is developed here based on the extended least action principle and assumptions of intrinsic angular momentum of an electron with random orientations. By incorporating appropriate relative entropy for the random orientations of intrinsic angular momentum in the extended least action principle, the theory recovers the quantum formulation of electron spin. The two-level quantization of spin measurement is a natural mathematical consequence. The formulation of measurement probability when a second Stern-Gerlach apparatus is rotated relative to the first Stern-Gerlach apparatus, and the Schrödinger-Pauli equation, are also derived successfully. Furthermore, we provide an intuitive physical model and formulation to explain the entanglement phenomenon between two electron spins. In this model, spin entanglement is the consequence of correlation between the random orientations of the intrinsic angular momenta of the two electrons. Since the orientation is an intrinsic local property of electron, the correlation of orientations can be preserved even when the two electrons are remotely separated. Such a correlation can be manifested without causal effect. Owing to this orientation correlation, the Bell-CHSH inequality is shown to be violated in a Bell test. A potential experiment is proposed to test the validity of the present theory. In a typical Bell test that confirms the violation of Bell-CHSH inequality, the present theory suggests that by adding a sufficiently large time delay before Bob’s measurement, the Bell-CHSH inequality can become non-violated. Limitations of the theory are also discussed to inspire further investigations on a more rigorous physical model of electron spin.
@@ I Introduction
With respect to spin entanglement between two electron spins, we will show that the root cause of entanglement is the correlation between the random variables, the orientations of the intrinsic angular momenta, of the two electrons. The correlation can be established during preparation. Crucially, since the orientation of the angular momentum is an intrinsic local property of the electron, the correlation can be maintained even though the two electrons are remotely separated and without interaction. Mathematically, we give an equivalent formulation of the four Bell states using probability density function instead of wave function. Based on the probability density function, Bell-CHSH inequality is proved to be violated in our theory due to the same root cause of entanglement. Since the correlation of the orientations of the angular momenta is preserved even when the two electrons are separated, the joint probability for the measurement outcome cannot be factorized into a product of two individual terms - a requirement for the Bell inequality to hold.
@@
The theory gives certain results that slightly different from standard quantum mechanics. Two potential experiments are proposed to test the difference and the validity of the theory. In the first experiment, the Stern-Gerlach apparatus is configured with a sufficiently weak gradient of the inhomogeneous magnetic field along the $z$-axis; We expect the electron detector screen will exhibit continuous distribution of displacements along the $z$-axis around two areas, instead of only two discrete lines. The second experiment is to modify a typical Bell test experiment that has already confirmed the violation of Bell-CSHS inequality. Here, we purposely let Alice and Bob perform their measurements at different times on each entangled pair of electrons. While Alice performs her measurement at time $t_{a}$, Bob purposely delays his measurement at time $t_{b}=t_{a}+\Delta t$. The present theory suggests that when $\Delta t$ is chosen properly, the Bell-CSHS inequality will not be violated.
@@
The rest of the paper is organized as follows. In Section II, we review the extended least action principle and its underlying assumptions. Detailed calculation showing how the basic quantum theory is derived from the principle is provided in Appendix A. Section III is devoted to developing the spin model, where the extended least action principle is incorporated with Tsallis relative entropy, resulting in a probability density function for the random orientation. The probability density function becomes two Dirac delta functions when the order of relative entropy approaches a limit, corresponding to the quantization of measurement outcomes. The theory for rotation of the spin is developed in this section as well. In Section IV, the Schrödinger-Pauli equation is derived by combining the effects of random translational fluctuations and random rotational fluctuations when applying the extended least action principle. Section V develops the theory for spin entanglement based on the correlations between the random orientations of intrinsic angular momenta, which allows us to prove that Bell inequality is violated with this spin theory. To test the validity of the theory, two potential experiments are proposed in Section VI.
There are clear limitations in the present theory, particularly the assumptions on the detailed physical model of electron spin, and the assumption that the order of relative entropy increases when the electron travels along an inhomogeneous magnetic field. A more rigorous theory must be developed to further justify these assumptions and beyond the scope of the present theory. The limitations of the present theory are discussed in Section VII.
@@ III.2 Discreteness of Spin Measurement
When an inhomogeneous field $B_{z}$ is applied, $m$ must be a number larger than zero. This means $m$ increases when the spatial gradient of $B_{z}$ is non-zero. We can assume further that $m$ monotonically increases as the electron travels along the inhomogeneous field $B_{z}$. The justification of this assumption requires a physical model of electron. We will provide an intuitive explanation of this assumption and discuss possible physical model in Section VII.2. Here we just call out this assumption explicitly.
@@ V.2 Singlet States
In the case that the initial orientations of angular momentum of electron $A$ are uniformly distributed for all $\theta\in[0,\pi]$, one has $\varrho(\Theta^{+}_{A})=\varrho(\Theta^{-}_{A})=1/2$, (62) becomes
@@
V.3 Other Bell States
On the other hand, if the correlations between the electron pair are such that $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ along the $z$-axis, and $\theta_{A}^{y}=\theta^{y}_{B}$ along the $y$-axis, we must also have
@@ V.3 Other Bell States
Similarly, we can give the equivalent descriptions of the other two Bell states in standard quantum mechanics. In the case that the initial orientations of angular momentum of electron $A$ are uniformly distributed for all $\theta\in[0,\pi]$, one has $\varrho(\Theta^{+}_{A})=\varrho(\Theta^{-}_{A})=1/2$, (60) becomes
the combination of (68) and (64) is equivalent to the Bell state
@@ VI Possible Experiments to Confirm the Spin Theory
The spin theory presented here so far recovers many of the results from standard quantum mechanics. However, when these conditions are not taken to the limits, the present theory will give some results different from standard quantum mechanics. In this section, we will provide two possible experiments to confirm the difference.
@@ VII.1 Conceptual Implications
We see the spin model and theory presented here gives an intuitive physical picture that may help to consolidate several quantum physics interpretations. In addition, our theory supports the idea that measurement outcome is context dependent. Measurement outcome for the spin of an electron depends on the measurement context, that is, the direction of magnetic field used in the Stern-Gerlach apparatus. The two level discreteness of measurement outcome is the result of a sufficiently strong gradient of magnetic field. For an electron with initially completely random orientation of intrinsic angular momentum, we can measure the spin along any direction but always get spin-up and spin-down with equal probability. There is no predefined spin property without specifying an experimental context.
@@ VII.2 Limitations
Secondly, assumption 4 suggests that when the interaction between the electron and an inhomogeneous magnetic field in a Stern-Gerlach apparatus takes place, the order of relative entropy monotonically increases, so that the direction of intrinsic angular momentum tends to align with the direction of the magnetic field. This assumption appears quite arbitrary, but can be intuitively explained as follows. The key is not to consider the electron as an idealized point particle. Instead it has non-zero size and is distributed with a spatial volume. We do not assume it is a rigid rotating body. Instead, it can be divided into many small mass elements $\delta m$. Each mass element is also a charge element $\delta e$, and experiences Larmor precession. Note that in Larmor precession, the angular frequency is $\omega=\delta eB_{z}/\delta m$. In a homogeneous magnetic field, $\omega$ is a constant for each charge element and independent of the angle $\theta$. In this case, the probability density function $p_{m}(\theta)$ should remain unchanged for the whole electron. But in an inhomogeneous field $B_{z}$, because the electron has spatial size, each charge element will experience a different magnitude of magnetic field. Consequently, different parts of the electron will have different angular frequencies with difference of $\Delta\omega$. The upper parts of the electron along the direction of $B_{z}$ tend to precess faster than the lower parts of the electron. Effectively, the orientation of overall intrinsic angular momentum is “pulled” closer to align with the $z$-axis. This corresponds to choosing a larger value of $m$. The spin theory based on stochastic mechanics Beyer seems to give similar results.
To estimate $\Delta\omega$, consider that in a typical Stern-Gerlach experiment, the gradient of magnetic field is of the order of $10^{3}$ Tesla/m. The size of electron is of order of $10^{-18}$m. Thus, the difference of magnetic field experienced at the upper part and lower part of the electron is about $\Delta B_{z}=10^{-15}$ Tesla. Then, the the Larmor precession frequency difference $\Delta\omega$ is at the order of $10^{-3}$Hz. On the other hand, typical Larmor frequency $\omega$ for an electron is $10^{10}$Hz. Thus, $\Delta\omega/\omega=10^{-13}$. This is an extremely small ratio. It raises the question on whether such a small ratio can cause the orientation of overall intrinsic angular momentum to tilt closer to align with the $z$-axis. A more rigorous theory to confirm this assumption is needed. Either considering the electron as an idealized charged point particle, or as a charge body with finite spatial size, seems insufficient as an accurate physical model for the electron. We speculate that using field theory could be a better framework to describe the physical model of electron, which is beyond the scope of this work but has been studied in Sebens .
@@ VII.3 Conclusions
From the extended least action principle, and with additional assumptions on the intrinsic angular momentum of an electron, particularly its random orientations, we are able to develop a theory that recovers the quantum formulation for electron spin. The key component of the theoretical framework here is the introduction of relative entropy for both translational and rotational random fluctuations. The two-level quantization of spin measurement outcomes is a mathematical consequence based on Assumption 4. We also obtain the same formulation as standard quantum mechanics when the direction of magnetic field of a second Stern-Gerlach apparatus is rotated with an angle from the direction of magnetic field of the first Stern-Gerlach apparatus. Recursively applying the extended least action principle, we have derived the Schrödinger-Pauli equation.
@@
When certain conditions are not taken to the limits, the present theory predicts results that are different from standard quantum mechanics. For instance, if the Stern-Gerlach apparatus is set up with a sufficiently weak gradient of the magnetic field along the $z$-axis, we expect the electron detector screen will exhibit continuous distribution of displacements along the $z$-axis, instead of only two discrete lines. Another more interesting experiment proposed is to add a delay before Bob’s measurement in the typical Bell test experiment, which could result in the non-violation of the Bell-CHSH inequality. The second experiment shows the effect of time dependency in measurements involving spin entanglement. If confirmed by experiment, such an effect can impose a limitation of the application of spin entanglement in quantum computing and quantum information.
The interplay of the extended least action principle and the information metrics for vacuum fluctuations prove to be very fruitful. We have shown the theoretical framework based on the extended least action principle has successfully recovered the non-relativistic quantum mechanics Yang2023 , the relativistic quantum scalar field theory Yang2024 , and in the paper, the quantum theory for electron spin. We believe the results in this paper are more interesting and because we can explain the entanglement phenomenon with an intuitive physical picture, and propose experiment to test potential new results different from standard quantum mechanics.
There are clear limitations in the present theory with the assumptions on the detailed physical model of electron spin. A more rigorous theory must be developed to further justify these assumptions. This paper is intended to be an incremental step for the investigation of spin theory using the information metrics for vacuum fluctuations.